OpenStudy (rainbow_dash):

Simplify. 5(a^2 + 5a + 6)/3(a^2 - 49) divided by 61(a + 3)/30(a + 7)

OpenStudy (rainbow_dash):

\[5(a^2+5a+6)/3(a^2-49) \div 61(a+3)/30(a+7)\]

OpenStudy (australopithecus):

fine

OpenStudy (australopithecus):

\[\frac{\frac{5(a^2+5a+6)}{3(a^2-49)}}{\frac{61(a+3)}{30(a+7)}}\] = \[\frac{(30(a+7))(5(a^2+5a+6)}{3(a^2-49)(61(a+3))}\]

OpenStudy (australopithecus):

\[\frac{(30(a+7))(5(a^2+5a+6))}{3(a^2-49)(61(a+3))} = \frac{50(a+7)(a^2+5a+6)}{61(a^2-49)(a+3)}\]

OpenStudy (australopithecus):

= \[\frac{50(a+7)(a+2)(a+3)}{61(a^2-49)(a+3)} = \frac{50(a+7)(a+2)}{61(a-7)(a+7)} = \frac{50(a+2)}{61(a-7)}\]

OpenStudy (rainbow_dash):

Fantastic. This helped a lot, thanks :)

OpenStudy (australopithecus):

is that thing pooping rainbows

OpenStudy (rainbow_dash):

It's actually a tail, but i can see how you think so. The picture is kind of small.

OpenStudy (australopithecus):

wow I'm learning

OpenStudy (rainbow_dash):

Every time you learn, you get a new wrinkle in your brain.

OpenStudy (australopithecus):

that is not how neurogenesis works

OpenStudy (rainbow_dash):

well i just got a new wrinkle....